MECHANISM OF THE THERMAL DECOMPOSITION OF FURAN

\underline{ANGAYLE VASILIOU}, G. BARNEY ELLISON, University of Colorado, Boulder, CO 80309-0215; MARK R. NIMLOS, Center for Renewable Chemical, Technologies & Materials, NREL, 1617 Cole Blvd., Golden, CO 80401; JOHN W. DAILY, Department of Mechanical Engineering, University of Colorado, Boulder, CO 80309-0427.

Both furan (C_{4}H_{4}O) and furfural (C_{4}H_{3}O-CHO) are important products in biomass pyrolysis. We have used a resistivelyheated SiC tubular reactor with a 30 microseconds residence time to study the thermal cracking of furan. The decomposition products are identified by two independent techniques: 118.2 nm VUV photoionization mass spectroscopy and infrared spectroscopy. We observe three different thermal dissociation channels leading to: a) CH_{3}CCH + CO b) HCCH + CH_{2}CO c) HCCCH_{2}.